# In-depth characterization of the Kepler-10 three-planet system with HARPS-N RVs and Kepler TTVs

### v1

@@ B.2 Apodized sine periodograms, white noise
The result from the two previous analyses show that there is a significant signal at 150.7 days, and hints of signals at 24 and 68 days, the latter being the most promising, but still not significant enough to be a clear detection. The question is then mainly to determine if the 150.7 days signal, which is clearly significant, can be attributed to a planet.
To address that question, we analyse the data in a different framework: the apodized sine periodogram (Hara et al. 2022a). This framework, proposed by Gregory (2016), is designed to determine if signals are present throughout the whole dataset, have a constant phase, amplitude and frequency. This approach is particularly useful when there is no clear model for the data. In practice, we fit apodized sine function,
$\mu_{K}(t,\omega,\tau,t_{0},A,B)=w(\tau,t_{0})(A\cos\omega t+B\sin\omega t).$ (11)
where $w(\tau,t_{0})$ is the apodization function, $t_{0}$ is its center and $\tau$ its temporal width. The resulting signal $\mu$ is then wavelet-like, quasi periodic from $\sim t_{0}-\tau$ to $\sim t_{0}+\tau$ and 0 elsewhere. We will here use Gaussian and box-shaped functions, that is
@@
where $\chi^{2}_{H}$ is the $\chi^{2}$ of the residuals after fitting the model $H$ (the base model, without sinusoidal function) and the model $K(t,\omega,\tau,t_{0})$, that is linearly fitting $A$ and $B$ for model Eq. 11 simultaneously with the base model. Then, for a given value of $\tau$, we represent $z$ as a function of $\omega$ maximised on $t_{0}$, that is for a given frequency and timescale, the best fitting model when the centre of the window varies.
@@
In Fig. 16 (a), we show the evolution of the phase and amplitude of that signal over the observations. While the phase is constant, the amplitude tends to vanish towards the end of the observations. We then multiplied the sinusoidal model of the signal by a so called apodization factor (Gregory 2016; Hara et al. 2022a), $\operatorname{e}^{(t-t_{0})^{2}/(2\tau^{2})}$ where $t$ is the time, and $t_{0}$ and $\tau$ are free parameters. The best fit is represented in Fig. 16 (b), where it seems that the amplitude of the signal vanishes at the end of the dataset. As said above, this is simply the best fitting timescale, but a purely sinusoidal signal ($\tau=10T_{\mathrm{obs}}$) is still compatible with the data. We performed the same analysis on the ancillary indicators and did not find trace of signals at $\sim 151$ d.
Figure 16: (a) Semi-Amplitude (red) and phase (green) of the 150 days signal averaged over a window of size $[t_{0}-\tau,t_{0}+\tau]$ where $\tau=T_{\mathrm{obs}}/3$, and $T_{\mathrm{obs}}$ is the total observation timespan. The solid lines represent the least square fit and shaded areas the $\pm$1 sigma intervals. Red and green dotted lines represent the mean value of the amplitude and phase, respectively. (b) In grey, HARPS-N data processed with YARARA-v2; in orange, a model of the form $\mathrm{e}^{-\frac{(t-t_{0})^{2}}{2\tau^{2}}}(A\cos\omega t+B\sin\omega t)$, where $\omega,A,B,\tau$ and $t_{0}$ are adjusted starting from $\omega=2\pi/150$ rad/day.
Appendix C Radial-velocity analyses.
@@
Table 11: Relative log Bayesian evidences, $\Delta\ln\mathcal{Z}$, as computed by PolyChord for non-GP and multi-dimensional (MD) GP models including different numbers of Keplerian terms, both with and without the inclusion of HIRES data.
@@
 [^15] Uncertainties correspond to the standard error on the mean evidence across three PolyChord runs, or to the highest individual run error provided by PolyChord (the greater of the two). For ease of interpretation, the highest evidence in a given column is normalized to zero; all other evidences in that column are defined relative to that zero point. Note that evidences cannot be compared meaningfully across setups (i)–(iv), as the data being modeled are different in each case. In each setup, a 3-planet model is strongly favored, with weak but insufficient statistical evidence in support of a 4-planet model.

### v2

@@ B.2 Apodized sine periodograms, white noise
The result from the two previous analyses show that there is a significant signal at 150.7 d, and hints of signals at 24 and 68 d, the latter being the most promising, but still not significant enough to be a clear detection. The question is then mainly to determine if the 150.7 d signal, which is clearly significant, can be attributed to a planet.
To address this question, we analyze the data in a different framework: the apodized sine periodogram (Hara et al. 2022a). This framework, proposed by Gregory (2016), is designed to determine if signals are present throughout the whole dataset, and have a constant phase, amplitude and frequency. This approach is particularly useful when there is no clear model for the data. In practice, we fit the apodized sine function
$\mu_{K}(t,\omega,\tau,t_{0},A,B)=w(\tau,t_{0})(A\cos\omega t+B\sin\omega t),$ (11)
where $w(\tau,t_{0})$ is the apodization function, $t_{0}$ is its center and $\tau$ its temporal width. The resulting signal $\mu$ is then wavelet-like, quasi periodic from $\sim t_{0}-\tau$ to $\sim t_{0}+\tau$ and 0 elsewhere. We here use Gaussian and box-shaped functions, that is
@@
where $\chi^{2}_{H}$ is the $\chi^{2}$ of the residuals after fitting the model $H$ (the base model, without sinusoidal function) and the model $K(t,\omega,\tau,t_{0})$, that is, linearly fitting $A$ and $B$ for model Eq. 11 simultaneously with the base model. Then, for a given value of $\tau$, we represent $z$ as a function of $\omega$ maximized on $t_{0}$, that is, for a given frequency and timescale, the best fitting model when the center of the window varies.
@@
In Fig. 15 (a) we show the evolution of the phase and amplitude of that signal over the observations. While the phase is constant, the amplitude tends to vanish towards the end of the observations. We then multiplied the sinusoidal model of the signal by a so called apodization factor (Gregory 2016; Hara et al. 2022a), $\operatorname{e}^{(t-t_{0})^{2}/(2\tau^{2})}$ where $t$ is the time, and $t_{0}$ and $\tau$ are free parameters. The best fit is represented in Fig. 15 (b), where it seems that the amplitude of the signal vanishes at the end of the dataset. As said above, this is simply the best fitting timescale, but a purely sinusoidal signal ($\tau=10\,T_{\mathrm{obs}}$) is still compatible with the data. We performed the same analysis on the ancillary indicators and did not find trace of signals at $\sim 151$ d.
Figure 15: (a) Semiamplitude (red) and phase (green) of the 150 d signal averaged over a window of size $[t_{0}-\tau,t_{0}+\tau]$ where $\tau=T_{\mathrm{obs}}/3$, and $T_{\mathrm{obs}}$ is the total observation timespan. The solid lines represent the least square fit, and the shaded areas the ±$\pm 1\sigma$ intervals. Red and green dotted lines represent the mean value of the amplitude and phase, respectively. (b) In grey, HARPS-N data processed with YARARA-v2; in orange, a model of the form $\mathrm{e}^{-\frac{(t-t_{0})^{2}}{2\tau^{2}}}(A\cos\omega t+B\sin\omega t)$, where $\omega,A,B,\tau$, and $t_{0}$ are adjusted starting from $\omega=2\pi/150$ rad/day.
Appendix C Radial-velocity analyses
@@
Table 11: Relative log Bayesian evidences, $\Delta\ln\mathcal{Z}$, as computed by PolyChord for non-GP and multidimensional (MD) GP models including different numbers of Keplerian terms, both with and without the inclusion of HIRES data.
@@
 [^15] Uncertainties correspond to the standard error on the mean evidence across three PolyChord runs, or to the highest individual run error provided by PolyChord (the greater of the two). For ease of interpretation, the highest evidence in a given column is normalized to zero; all other evidences in that column are defined relative to that zero point. Note that evidences cannot be compared meaningfully across setups (i)–(iv), as the data being modeled are different in each case. In each setup, a three-planet model is strongly favored, with weak but insufficient statistical evidence in support of a four-planet model.